OpenStudy (anonymous):

correct way to write this mixed number as a decimal number 8 4/10

OpenStudy (austinl):

\[8 \frac{4}{10}\] = \[\frac{84}{10}\] 84 divided by 10 is 8.4

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

How do you read 4/10? Four tenths. Four tenths is 0.4 8 4/10 is eight and four tenths, so it's 8 + 4/10 = 8 + 0.4 = 8.4

OpenStudy (anonymous):

FIRST WRITE THE WHOLE NUMBER IN THE LEFT THEN COVERT THE FRACTION INTO DECIMAL N WRITE TO THE RIGHT OF DECIMAL I.E. first wrie 8 to the left of decimal then convert 4/10 to decimal which is 0.4 nw add these u l get 8.4